# On denominators of consecutive $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions

### v1

@@ Abstract.
The sequences of $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions were defined in our previous work as ${\mathscr{S}}_{Q}:=\{a/q\in{\mathbb{Q}}\cap(0,1]:q+a+\bar{a}\leq Q\}$, where $\bar{a}$ is the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. Here, we prove that the set of $Q$-scaled denominators of consecutive fractions in ${\mathscr{S}}_{Q}$ is dense in the region ${\mathcal{V}}:=\{(x,y)\in[0,1]^{2}:\max\{(1-3x)/2,2x-1\}\leq y\leq\max\{x,1-x\}\}$, and provide a formula for their distribution in $\mathcal{V}$} as $Q\rightarrow\infty$.
@@ 1. Introduction
$\mathcal{F}_{Q}:=\left\{\frac{d}{b}:d,b\in\mathbb{N},\ 1\leq d\leq b\leq Q,\ (d,b)=1\right\}.$
@@
where $\bar{a}$ denotes the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. In particular, this shows that $\mathcal{F}_{Q}\subseteq\mathscr{S}_{3Q}$. Note also that $h(a/q)\neq h(1-a/q)$, but some sort of $(x\mapsto 1-x)$-symmetry of the height $h$ is present in the identity
@@
The sets $\mathscr{S}_{Q}$ exhibit interesting arithmetic properties, resembling in part those of $\mathcal{F}_{Q}$. Firstly, their union gives precisely the set ${\mathbb{Q}}\cap(0,1]$. Secondly, as shown in the appendix to [3], the elements of each set $\mathscr{S}_{Q}^{*}:=\{0\}\cup\mathscr{S}_{Q}$ define a modular partition of the interval $[0,1]$. This means that if $0/1<a_{1}/q_{1}=1/(Q-2)<\cdots<a_{\#\mathscr{S}_{Q}}/b_{\#\mathscr{S}_{Q}}=1/1$ denotes the elements of $\{0\}\cup\mathscr{S}_{Q}$, then $a_{i+1}q_{i}-a_{i}q_{i+1}=1$ for any $i=1,\ldots,\#\mathscr{S}_{Q}-1$.
In sharp contrast with the sequence $(\mathcal{F}_{Q})_{Q}$, the elements of $(\mathscr{S}_{Q})_{Q}$ are no longer uniformly distributed as $Q\rightarrow\infty$. Furthermore, it was proved in [3, Theorem 1] that
$\big{(}\mathscr{S}_{Q}\cap[0,\beta]\big{)}=\frac{Q^{2}}{2\zeta(2)}\log\bigg{(}\frac{2+2\beta}{2+\beta}\bigg{)}+O_{\varepsilon}(Q^{3/2+\varepsilon}),\quad\forall\beta\in[0,1].$
@@
One of the fundamental features of the Farey sequence $(\mathcal{F}_{Q})_{Q}$ is that the elements of the set $\mathcal{F}^{*}_{Q}:=\{0\}\cup\mathcal{F}_{Q}$ form a modular partition of the interval $[0,1]$. In particular, the elements of $\mathcal{F}^{*}_{Q}$ are determined by their denominators. The elements of the set
$\mathcal{D}_{Q}:=\big{\{}(q_{1},q_{2}):\exists\,a_{1},a_{2}\in\mathbb{Z},\ a_{1}/q_{1}<a_{2}/q_{2}\text{ consecutive elements in }\mathcal{F}^{*}_{Q}\big{\}}$
are in one-to-one correspondence with the set of primitive lattice points in the triangle $Q\mathcal{T}$, where $\mathcal{T}$ denotes the Farey triangle $\{(x,y)\in[0,1]^{2}:x+y>1\}$. As a result, the scaled set $\mathcal{D}:=\cup_{Q}Q^{-1}\mathcal{D}_{Q}$ is dense in the closed triangle $\overline{\mathcal{T}}$. It was first noticed by Kargaev and Zhigljavsky [13] (see also [6, Lemma 2] for a quick effective proof) that the $Q$-scaled sets $Q^{-1}\mathcal{D}_{Q}$ are uniformly distributed in $\overline{\mathcal{T}}$, in the sense that, for every rectangular box $\mathcal{R}\subseteq\overline{\mathcal{T}}$,
$\lim\limits_{Q}\frac{\#(\mathcal{D}_{Q}\cap Q\mathcal{R})}{\#\mathcal{D}_{Q}}=\operatorname{Area}(\mathcal{R}).$ (2)
As noticed in [6], the shifting of consecutive elements in $\mathcal{D}_{Q}$ is conveniently described by iterating the map $T:\mathcal{T}\rightarrow\mathcal{T}$,
@@
which satisfies, for every $Q\geq 1$,
@@
$\begin{split}&\mathcal{V}_{1}:=\bigg{\{}(x,y)\in\bigg{[}\frac{1}{2},1\bigg{]}\times\bigg{[}\frac{1}{3},1\bigg{]}:\max\{1-x,2x-1\}\leq y\leq x\bigg{\}},\\ &\mathcal{V}_{2}:=\bigg{\{}(x,y)\in\bigg{[}0,\frac{1}{2}\bigg{]}\times\bigg{[}\frac{1}{5},1\bigg{]}:\max\bigg{\{}\frac{1-3x}{2},x\bigg{\}}\leq y\leq 1-x\bigg{\}},\\ &\mathcal{V}_{3}:=\bigg{\{}(x,y)\in\bigg{[}\frac{1}{5},\frac{2}{3}\bigg{]}\times\bigg{[}0,\frac{1}{2}\bigg{]}:\max\bigg{\{}\frac{1-3x}{2},0,2x-1\bigg{\}}\leq y\leq\min\{x,1-x\}\bigg{\}}.\end{split}$ (3)
@@
$\lim\limits_{Q}\frac{\#(\mathscr{D}_{Q}\cap Q\mathcal{R})}{\#\mathscr{D}_{Q}}=\iint_{\mathcal{R}}H_{i}(x,y)\,dxdy,$
@@
The region $\mathcal{V}$ can be partitioned into subregions $\mathcal{W}_{1},\mathcal{W}_{2},\mathcal{W}_{3},\ldots$, with $\mathcal{W}_{r}$ representing the closure of the scaled set $\cup_{Q}Q^{-1}\mathscr{D}_{Q}^{(r)}$, where $\mathscr{D}_{Q}^{(r)}$ denotes the set of pairs of denominators of consecutive elements of $\mathscr{S}_{Q}$ with exactly r1r-1 elements of $\mathcal{F}_{Q}$ in between (see the representation in the right side of Figure 2). In addition, every $\mathcal{W}_{r}$, $r\geq 3$, can be further partitioned into r1r-1 (possibly empty) polygons $\mathcal{W}_{r,i}$, with $i$ being the (unique) place where the mediant of consecutive denominators is being attained, that is $q_{i}=q_{1}+q_{r+1}$. The boundaries of the polygons $\mathcal{W}_{r,i}$ are computable, as illustrated in Section 4 for $r\leq 5$. See also Figures 1 and 2.
If the rational numbers $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ satisfy $a_{2}q_{1}-a_{1}q_{2}=1=a_{3}q_{2}-a_{2}q_{3}$, then
@@
Furthermore, when $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ are consecutive in $\mathcal{F}_{Q}$, we have
@@
One nice feature of the index, proved in [10, Theorem 1] through an elementary mediant argument, is the identity
@@ 2. The inclusion $\mathscr{D}\subseteq\mathcal{V}$
Let $r\geq 1$. Suppose that $a_{1}/q_{1}<\cdots<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathcal{F}_{Q}$ with $a_{1}/q_{1},a_{r+1}/q_{r+1}\in\mathscr{S}_{Q}$ and $a_{2}/q_{2},\ldots,a_{r}/q_{r}\notin\mathscr{S}_{Q}$ (in other words $a_{1}/q_{1}<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathscr{S}_{Q}$). Denote by $\mathscr{D}_{Q}^{(r)}$ the set of denominator pairs $(q_{1},q_{r+1})$ of consecutive elements of $\mathscr{S}_{Q}$, with $q_{1}$ and $q_{r+1}$ as above.
@@
As in Section 3.1 of [3] and employing the notation thereby, in this case we have $q_{2}<q_{1}$, so $q_{1}\in[Q/2,Q]$, and also $3q_{1}-Q-q_{2}\leq\bar{q}_{2}<q_{1}$, leading to $q_{2}>2q_{1}-Q$. We also employ the trivial inequality $q_{2}>Q-q_{1}$. ∎
@@
$a_{r+1}=K^{\prime}a_{2}-K^{\prime\prime}a_{1}.$
@@
Upon Lemma 6, $q_{1}+q_{r+1}=q_{i}$ for some $i\in\{2,\ldots,r\}$. We have
@@
The equality (7) yields $q_{r}=q_{1}+\nu q_{r+1}$. It is plain that $K\geq 1$ and $\nu\geq 1$ in (7). Indeed, if $K\leq-1$, then $Q\geq q_{r+1}=q_{2}-Kq_{1}\geq q_{2}+q_{1}>Q$, a contradiction, while $K=0$ yields $(q_{r+1},q_{r})=(q_{2},q_{1})$, a contradiction as $r>1$. Hence $K\geq 1$. Next, if $\nu\leq-1$, then $q_{r}\leq q_{1}-q_{r+1}$, so $Q<q_{r}+q_{r+1}\leq q_{1}\leq Q$, a contradiction. Finally, if $\nu=0$, then $q_{r}=q_{1}$ and $Q<q_{r}+q_{r+1}=q_{1}+q_{2}-Kq_{1}\leq q_{2}\leq Q$, a contradiction. Hence $\nu\geq 1$ and we infer $q_{1}+q_{r+1}\leq q_{1}+\nu q_{r+1}=q_{r}\leq Q$.
It was observed in Sections 3.2 and 3.3 of [3] that when $r\geq 2$ we have $q_{1}<q_{2}$. Next, we prove the inequality $q_{r+1}-2q_{1}+Q\geq 0$. This is obsolete when $K\geq 2$, as in this case we can use $q_{r+1}=q_{2}-Kq_{1}$ to infer $q_{1}\leq q_{2}/K\leq Q/2$. When $K=1$, we have $q_{r+1}-2q_{1}=q_{2}-3q_{1}$. Now, the first inequality in [3, (16)], where $q=q_{2}>\alpha=q_{1}$, yields $3q_{1}-q_{2}\leq(2+[q_{2}/q_{1}])q_{1}-q_{2}+a_{1}\leq Q$, which leads to $q_{2}-3q_{1}+Q\geq 0$.
Finally, the inequality $3q_{1}+2q_{r+1}\geq Q$ was proved in Lemma 5. ∎
3. The inclusion $\mathcal{V}\subseteq\overline{\mathscr{D}}$
We now show that $\mathcal{V}$ cannot be further restricted; that is that $\mathcal{V}\subseteq\overline{\mathscr{D}}$, the closure of the set $\mathscr{D}$ defined in Theorem 1. In particular, we prove the following density result.
Proposition 9.
For any $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{1}\cup\mathring{\mathcal{V}}_{2}\cup\mathring{\mathcal{V}}_{3}$ and $\varepsilon_{1},\varepsilon_{2}>0$, there are $Q\geq 3$ and $q_{1},q_{2}$ such that
$(q_{1},q_{2})\in\mathscr{D}_{Q}\cap Q\big{(}[x_{0}-\varepsilon_{1},x_{0}+\varepsilon_{1}]\times[y_{0}-\varepsilon_{2},y_{0}+\varepsilon_{2}]\big{)}.$
We simply denote $\mathcal{R}:=[x_{0}-\varepsilon_{1},x_{0}+\varepsilon_{1}]\times[y_{0}-\varepsilon_{2},y_{0}+\varepsilon_{2}]=:I_{1}\times I_{2}$.
The statement of Theorem 2 will be proved in parallel.
We start off with the easiest case $x_{0}+y_{0}>1$ and assume, without loss of generality, that $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{1}$. Here we can look for $q_{1},q_{2}$ denominators of consecutive elements in $\mathcal{F}_{Q}$, so that $q_{1}+q_{2}>Q$. The considerations from Section 3.1 of [3] (see especially Corollary 6 thereby) yield

### v2

@@ Abstract.
The sequence $({\mathscr{S}}_{Q})_{Q}$ of $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions was defined in our previous work by ${\mathscr{S}}_{Q}:=\{a/q\in{\mathbb{Q}}\cap(0,1]:q+a+\bar{a}\leq Q\}$, where $\bar{a}$ is the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. Here, we prove that the set of $Q$-scaled denominators of consecutive fractions in ${\mathscr{S}}_{Q}$ is dense in the region ${\mathcal{V}}:=\{(x,y)\in[0,1]^{2}:\max\{(1-3x)/2,2x-1\}\leq y\leq\max\{x,1-x\}\}$, and provide a formula for their distribution in $\mathcal{V}$} as $Q\rightarrow\infty$.
@@ 1. Introduction
$\mathcal{F}_{Q}:=\left\{\frac{d}{b}:d,b\in\mathbb{N},\ 1\leq d\leq b\leq Q,\ \gcd(d,b)=1\right\}.$
@@
where $\bar{a}$ denotes the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. In particular, this shows that $\mathcal{F}_{Q}\subseteq\mathscr{S}_{3Q}\subseteq\mathcal{F}_{3Q}$, and so
$\bigcup_{Q\geq 3}\mathscr{S}_{Q}=\bigcup_{Q\geq 1}\mathcal{F}_{Q}={\mathbb{Q}}\cap(0,1].$ (1)
Note also that $h(a/q)\neq h(1-a/q)$, but some sort of $(x\mapsto 1-x)$-symmetry of $h$ is present in the identity
@@
To compare the sets $\mathscr{S}_{Q}$ and $\mathcal{F}_{Q}$, look at
$\mathscr{S}_{10}=\bigg\{\frac{1}{8},\frac{1}{7},\frac{1}{6},\frac{1}{5},\frac{1}{4},\frac{1}{3},\frac{2}{5},\frac{1}{2},\frac{3}{5},\frac{2}{3},\frac{3}{4},\frac{1}{1}\bigg\}$
and
$\mathcal{F}_{10}\setminus\mathscr{S}_{10}=\bigg\{\frac{1}{10},\frac{1}{9},\frac{2}{9},\frac{2}{7},\frac{3}{10},\frac{3}{8},\frac{3}{7},\frac{4}{9},\frac{5}{9},\frac{4}{7},\frac{5}{8},\frac{7}{10},\frac{5}{7},\frac{7}{9},\frac{4}{5},\frac{5}{6},\frac{6}{7},\frac{7}{8},\frac{8}{9},\frac{9}{10}\bigg\}.$
Besides property (1), the sets $\mathscr{S}_{Q}$ exhibit another interesting arithmetic property resembling $\mathcal{F}_{Q}$. As shown in the appendix to [3], the elements of each set $\mathscr{S}_{Q}^{*}:=\{0\}\cup\mathscr{S}_{Q}$ define a unimodular partition of the interval $[0,1]$. This means that if $0/1<a_{1}/q_{1}=1/(Q-2)<\cdots<a_{\#\mathscr{S}_{Q}}/b_{\#\mathscr{S}_{Q}}=1/1$ denote the elements of $\mathscr{S}^{*}_{Q}$, then $a_{i+1}q_{i}-a_{i}q_{i+1}=1$ for any $i=1,\ldots,\#\mathscr{S}_{Q}-1$.
However, in sharp contrast with $(\mathcal{F}_{Q})_{Q}$, the elements of $(\mathscr{S}_{Q})_{Q}$ are no longer uniformly distributed as $Q\rightarrow\infty$. Furthermore, it was proved in [3, Theorem 1] that
$\big(\mathscr{S}_{Q}\cap[0,\beta]\big)=\frac{Q^{2}}{2\zeta(2)}\log\bigg(\frac{2+2\beta}{2+\beta}\bigg)+O_{\varepsilon}(Q^{3/2+\varepsilon}),\quad\forall\beta\in[0,1].$ (2)
@@
It is well-known that each set $\mathcal{F}^{*}_{Q}:=\{0\}\cup\mathcal{F}_{Q}$ defines a unimodular partition of the interval $[0,1]$. In particular, the elements of $\mathcal{F}^{*}_{Q}$ are determined by their denominators only. The elements of the set
$\mathcal{D}_{Q}:=\big\{(q_{1},q_{2})\in\mathbb{N}^{2}:\exists\,a_{1},a_{2}\in\mathbb{Z},\ a_{1}/q_{1}<a_{2}/q_{2}\text{ consecutive elements in }\mathcal{F}^{*}_{Q}\big\}$
are in one-to-one correspondence with the set of primitive lattice points in the triangle $Q\mathcal{T}$, where $\mathcal{T}$ denotes the Farey triangle $\{(x,y)\in[0,1]^{2}:x+y>1\}$. As a result, the scaled set $\mathcal{D}:=\cup_{Q}Q^{-1}\mathcal{D}_{Q}$ is dense in the closed triangle $\overline{\mathcal{T}}$. It was first noticed by Kargaev and Zhigljavsky [14] (see also [7, Lemma 2] for a quick effective proof) that the $Q$-scaled sets $Q^{-1}\mathcal{D}_{Q}$ are uniformly distributed in $\overline{\mathcal{T}}$, in the sense that, for every rectangular box $\mathcal{R}\subseteq\overline{\mathcal{T}}$,
$\lim\limits_{Q}\frac{\#(\mathcal{D}_{Q}\cap Q\mathcal{R})}{\#\mathcal{D}_{Q}}=2\operatorname{Area}(\mathcal{R}).$ (4)
As noticed in [7], the shifting of denominators of consecutive elements in $\mathcal{F}_{Q}^{*}$ is conveniently described by iterating the map $T:\mathcal{T}\rightarrow\mathcal{T}$,
@@
which is invertible, area-preserving, and satisfies, for every $Q\geq 1$,
@@
We will first show the inclusion $\mathscr{D}\subseteq\mathcal{V}$, then prove Theorem 2, a stronger result which provides an asymptotic formula for the cardinality of the set $\mathscr{D}_{Q}\cap Q\mathcal{R}$ for rectangular boxes $\mathcal{R}$ inside $\mathcal{V}$. This will show in particular that $Q^{-1}\mathscr{D}_{Q}\cap\mathcal{R}\neq\emptyset$ for large enough $Q$ (depending on $\mathcal{R}$).
@@
$\begin{split}&\mathcal{V}_{1}:=\bigg\{(x,y)\in\bigg[\frac{1}{2},1\bigg]\times\bigg[\frac{1}{3},1\bigg]:\max\{1-x,2x-1\}\leq y\leq x\bigg\},\\ &\mathcal{V}_{2}:=\bigg\{(x,y)\in\bigg[\frac{1}{5},\frac{2}{3}\bigg]\times\bigg[0,\frac{1}{2}\bigg]:\max\bigg\{\frac{1-3x}{2},0,2x-1\bigg\}\leq y\leq\min\{x,1-x\}\bigg\},\\ &\mathcal{V}_{3}:=\bigg\{(x,y)\in\bigg[0,\frac{1}{2}\bigg]\times\bigg[\frac{1}{5},1\bigg]:\max\bigg\{\frac{1-3x}{2},x\bigg\}\leq y\leq 1-x\bigg\}.\end{split}$ (5)
@@
$\lim\limits_{Q}\frac{\#(\mathscr{D}_{Q}\cap Q\mathcal{R})}{\#\mathscr{D}_{Q}}=\frac{2}{\log(4/3)}\iint_{\mathcal{R}}H_{i}(x,y)\,dxdy,$
@@
The region $\mathcal{V}$ can be partitioned into subregions $\mathcal{W}_{1},\mathcal{W}_{2},\mathcal{W}_{3},\ldots$, with $\mathcal{W}_{r}$ representing the closure of the scaled set $\cup_{Q}Q^{-1}\mathscr{D}_{Q}^{(r)}$, where $\mathscr{D}_{Q}^{(r)}$ denotes the set of pairs of denominators of consecutive elements of $\mathscr{S}_{Q}$ with exactly r1r-1 elements of $\mathcal{F}_{Q}$ in between (see the representation on the right-hand side of Figure 2). In addition, every $\mathcal{W}_{r}$, $r\geq 3$, can be further partitioned into r1r-1 (possibly empty) polygons $\mathcal{W}_{r,i}$, with $i\in\{2,\ldots,r\}$ being the (unique) place where the mediant of consecutive denominators is being attained, that is $q_{i}=q_{1}+q_{r+1}$. The boundaries of the polygons $\mathcal{W}_{r,i}$ are computable, as illustrated in Section 4 for $r\leq 5$. See also Figures 1 and 2.
If the non-negative rational numbers $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ satisfy $a_{2}q_{1}-a_{1}q_{2}=1=a_{3}q_{2}-a_{2}q_{3}$, then
@@
Furthermore, when $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ are consecutive in $\mathcal{F}_{Q}^{*}$, we have
@@
One nice feature of the index, proved in [11, Theorem 1] through an elementary mediant argument, is the identity
@@ 2. The inclusion $\mathscr{D}\subseteq\mathcal{V}$
Let $r\geq 1$. Suppose $a_{1}/q_{1}<\cdots<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathcal{F}_{Q}$, $a_{1}/q_{1},a_{r+1}/q_{r+1}\in\mathscr{S}_{Q}$ and $a_{2}/q_{2},\ldots,a_{r}/q_{r}\notin\mathscr{S}_{Q}$ (in other words $a_{1}/q_{1}<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathscr{S}_{Q}$). Denote by $\mathscr{D}_{Q}^{(r)}$ the set of denominator pairs $(q_{1},q_{r+1})$ of consecutive elements of $\mathscr{S}_{Q}$, with $q_{1}$ and $q_{r+1}$ as above.
@@
As in Section 3.1 of [3], in this case we have $q_{2}<q_{1}$, so $q_{1}\in[Q/2,Q]$, and also $3q_{1}-Q-q_{2}\leq\bar{q}_{2}<q_{1}$ with $\bar{q}_{2}$ denoting the multiplicative inverse of $q_{2}\left(\mathrm{mod\;}q_{1}\right)$ in $1,q_{1})$, leading to $q_{2}>2q_{1}-Q$. We also employ the trivial inequality $q_{2}>Q-q_{1}$. ∎
@@
$q_{r+1}=K^{\prime}q_{2}-K^{\prime\prime}q_{1}\quad\text{and}\quad a_{r+1}=K^{\prime}a_{2}-K^{\prime\prime}a_{1}.$
@@
As $a_{1}/q_{1}<\cdots<a_{r}/q_{r}$ are consecutive in $\mathcal{F}_{Q}$, Lemma 6 yields $q_{1}+q_{r+1}=q_{i}$ for some $i\in\{2,\ldots,r\}$. We have
@@
Lemmas 6 and 7 provide $q_{1}+q_{r+1}\leq Q$ and $3q_{1}+2q_{r+1}\geq Q$, so it remains to prove the inequality $q_{r+1}-2q_{1}+Q\geq 0$.
With $K$ as in equality (9), the last inequality is obsolete when $K\geq 2$. Indeed, it was observed in Sections 3.2 and 3.3 of [3] that when $r\geq 2$ we have $q_{1}<q_{2}$. and we can use $q_{r+1}=q_{2}-Kq_{1}$ to infer $q_{1}\leq q_{2}/K\leq Q/2$. When $K=1$, we have $q_{r+1}-2q_{1}=q_{2}-3q_{1}$. Now, the first inequality in [3, (16)], where $q=q_{2}>\alpha=q_{1}$, yields $3q_{1}-q_{2}\leq(2+[q_{2}/q_{1}])q_{1}-q_{2}+a_{1}\leq Q$, which leads to $q_{2}-3q_{1}+Q\geq 0$. Finally, we notice that $K\leq 0$ cannot occur, as $K\leq-1$ would imply $Q\geq q_{r+1}=q_{2}-Kq_{1}\geq q_{2}+q_{1}>Q$, a contradiction, while $K=0$ yields $q_{r+1}=q_{2}$ and $q_{r}=q_{1}$, a contradiction as $r>1$. ∎
3. The inclusion $\mathcal{V}\subseteq\overline{\mathscr{D}}$ and the proof of Theorem 2
We start off the proof of Theorem 2 by fixing $(x_{0},y_{0})\in\mathcal{R}$ and $\varepsilon_{1},\varepsilon_{2}>0$ such that
$\mathcal{R}:=[x_{0}-\varepsilon_{1},x_{0}+\varepsilon_{1}]\times[y_{0}-\varepsilon_{2},y_{0}+\varepsilon_{2}]=:I_{1}\times I_{2}\subseteq\mathring{\mathcal{V}}_{1}\cup\mathring{\mathcal{V}}_{2}\cup\mathring{\mathcal{V}}_{3}.$
We will count the number of integer pairs $(q_{1},q_{2})\in\mathscr{D}_{Q}\cap Q\mathcal{R}$. The situation where $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{1}$, so $x_{0}+y_{0}>1$, $x_{0}>y_{0}$ and $1+y_{0}-2x_{0}>0$, is considered first. Here, we can look for $q_{1},q_{2}$ with $q_{1}+q_{2}>Q>2q_{1}-q_{2}$ and $q_{1}>q_{2}$. Hence, they are denominators of consecutive elements in $\mathcal{F}_{Q}$ and the considerations from Section 3.1 of [3] (see especially Corollary 6 thereby) yield